# SAT: Dynamic Spatial Aptitude Training for Multimodal Language Models

### v2

@@ 3.3 Finetuning MLMs with SAT
We provide the MLM with the required image(s) and the question and ask it to choose between two answers—a correct answer and a distractor answer, following existing benchmark standards (Fu et al., 2024b; Cai et al., 2024). The distractor answer is generated by switching the correct answer to the opposite/distractor word (e.g., left $\rightarrow$ right, or “did not move”). To prevent catastrophic forgetting, we mix in some of the original pre-training data during tuning from the LLaVA Instruct-tune dataset (Liu et al., 2024). More tuning and prompt details are in the appendix.
@@ 4 Experiments
Setup. We use two popular open-source MLMs: LLaVA-1.5-13B (Liu et al., 2024) and LLaVA-Video-7B (Zhang et al., 2024c) for our experiments. We report the standard accuracy metric used for question-answering evaluations by checking if the predicted answer matches the GT answer. We notice better performance for off-the-shelf MLMs when we provide the options in text (e.g., Choose between right or left) as opposed to option numbers (e.g., Choose between option A or B). Hence, we sometimes report higher performances for the baselines than those reported by the original papers. Below, we present the key experimental questions we wish to ask.
@@
CVBench (Tong et al., 2024) BLINK (Fu et al., 2024b)
@@
Table 4: SAT also improves spatial performance on videos, VSI-Bench (Yang et al., 2024a)
@@
Table 5: Performance on some traditional spatial benchmarks- VSR (Liu et al., 2023), GQA-Sp (Hudson & Manning, 2019b)- and other VQA benchmarks- GQA (Hudson & Manning, 2019b), OK-VQA (Marino et al., 2019), VQAv2 (Goyal et al., 2017)- not focused on spatial relationships. Our spatial tuning remembers pre-training commonsense on other capabilities while improving spatial reasoning.
@@ 4.1 Can SAT improve spatial reasoning in MLMs?
To evaluate the effect of SAT training on spatial performance, we use 4 existing static spatial benchmarks—CVBench (Tong et al., 2024), BLINK (spatial subsets) (Fu et al., 2024b), Visual Spatial Relations (VSR) (Liu et al., 2023), and GQA-Spatial (Hudson & Manning, 2019a). For dynamic reasoning, we use our SAT test sets on both real and synthetic images. Since SAT real is small, we only report the overall average and report individual split performance on our synthetic test set for a more detailed analysis. Finally, we evaluate the effect of SAT tuning on a recent long-video spatial benchmark, VSI-Bench (Yang et al., 2024a), by tuning LLaVA-Video-7B (Zhang et al., 2024c).
Open MLMs struggle on spatial reasoning. While proprietary MLMs perform better on static reasoning, they still struggle with dynamic reasoning. In Tables 3 and 2, we see that open-source MLMs (rows f and h) struggle on both static relationships and dynamic reasoning. Larger closed-source models like GPT4-o (Achiam et al., 2023), Gemini-1.5-pro (Team et al., 2023) and spatially-tuned Robopoint (Yuan et al., 2024b) are better at static reasoning (in Table 3 row b, d), but still struggle on dynamic QAs (Table 2 rows a-e). On SAT synthetic, aiming at the goal is generally easier for spatially stronger models like RoboPoint and GPT4-o since it also mostly requires judging object position.
Training on SAT significantly improves static spatial reasoning on real images for MLMs. As shown in Table 3, tuning on SAT improves performance on static spatial questions for both LLaVA-1.5 and LLaVA-Video models by significant amounts (rows f vs g and h vs i). Comparing rows a-e, we observe that SAT tuning makes both LLaVA models match/outperform some closed-source models on zero-shot performance on CVBench and BLINK. Compared to an existing spatially-tuned baseline, we outperform RoboPoint-13B (Yuan et al., 2024b) (row e). This shows promise that SAT instruction-tuning may push performance further for these stronger models.
@@
Camera movement and out-of-domain relations are hard to improve BLINK SpRel has abstract relationships (especially with people) not present in our synthetic data (e.g., “looking away”, “surrounding”). Hence, in Table 3, gains are modest on SpRel (row f vs g), with LLaVA-Video (row h vs i) showing no improvement. Camera movements are also hard (BLINK MV). While more investigation is needed, we believe this may be due to translation invariance in ViT (Dosovitskiy et al., 2021) leading to minimal feature changes between images with subtle camera movement.
SAT synthetic is also a useful diagnostic test set The lack of improvement on object relations (SpRel and 2DRel) in LLaVA-Video is also reflected by a lesser gain on the GoalAim (row h vs i, Table 2) on SAT synthetic test, where the model needs to judge object position. Similar for SAT EgoM (Table 2) and BLINK MV (Table 3) dealing with camera movements-LLaVA-Video improves more than LlaVA on both. Hence, despite being synthetic, SAT synthetic is also a useful diagnostic test set.
Our SAT-tuned model maintains its non-spatial reasoning capabilities. We run an evaluation on some standard VQA benchmarks—namely, GQA (Hudson & Manning, 2019a), VQAv2 (Goyal et al., 2017), and OK-VQA (Marino et al., 2019) (9K image-QA pairs, random 3K from each). We maintain overall performance on them compared to the off-the-shelf LLaVA (Liu et al., 2024) as shown in Table 5. This suggests that we remember pre-trained vision-language commonsense while adding stronger spatial capabilities.
Training on SAT improves spatial reasoning in long videos. Even though SAT only has up to 2 frames in its training data, we evaluate its impact on longer videos on Yang et al. (2024a). The results are shown in Table 4. We see that SAT training surprisingly helps significantly on route planning capabilities (column Rt. Plan) while also improving other spatial splits. This shows promise that SAT tuning can benefit embodied AI applications.
Table 6: Table showing the effectiveness of SAT over existing sources of spatial data. Dynamic spatial reasoning data improves performance over just static spatial relationship data.
@@ 4.2 What are effective types of spatial training data?
Sources of instruction tuning Since contemporary work on spatial understanding uses pseudo annotations (Chen et al., 2024; Cheng et al., 2024), we wish to compare to such baselines as well. Since their datasets are not easily available yet, we reproduce a similar instruction-tuning set using pseudo annotations to create spatial QAs. We keep formats the same as SAT to ensure the performance differences are attributed to the source of spatial information and not the question formats.
  • GQAPseudo: We create SAT-Static like instruction-tuning data by inferring the depth (Yang et al., 2024b) on images from GQA (Hudson & Manning, 2019a) and VisualGenome (Krishna et al., 2017) and the 2D bounding box annotations. We filter out potential incorrect relationships using simple heuristics (more details in the appendix). We create $225$K static spatial image-QA tuples. Generating pseudo-annotated dynamic QAs is not possible since we cannot easily infer the result of actions on real images, hence, we only generate static QAs.
  • VSR/25VRD: We use spatial relationship datasets like VSR (Liu et al., 2023) and 2.5VRD (Su et al., 2021) to produce more real spatial QAs. These datasets contain relationships such as “touching”, or “behind”, allowing us to formulate QAs such as “Is the cat touching the sofa?” We combine VSR and 2.5VRD to create $200$K image-QA tuples.
@@
Figure 4: Ablations with training data size. We see that dynamic data scales better with data, whereas static tends to saturate.
SAT simulated data outperforms pseudo-annotations or human annotations. In Table 6, we see that pseudo-annotations can improve performance on certain splits like Count and 2DRel on CVBench (row b). However, noise in annotations (e.g., bounding boxes often not accounting for the entire object) leads to errors in judging depth of the entire object relative to other objects (e.g., a part of the annotated “lake” may be in front of a “tree”, but the majority of it may lie behind it). Following more complex careful curation similar to Cheng et al. (2024) may improve this performance, which we leave to future work since their data is not yet easily available. Human-annotated spatial relations (using Su et al. (2021); Liu et al. (2023)) tend to perform well in-domain on VSR test. While the higher diversity of human-annotated relations also helps prevent forgetting on BLINK SpRel (row c vs d), they are finite and not easily composable to generate varied 3D and dynamic data. Hence, we observe no improvements on SAT dynamic and minimal gains in 3D perception splits.
@@
SAT training with dynamic data scales positively with more data Since SAT is generated procedurally, we wish to explore whether spatial performance of MLMs scale with more data. Hence, we vary the amounts of training data (train for 1 epoch). As shown in Figure 4, we see that while performance tends to saturate with static-only training (likely due to overfitting to simpler relationships), the performance scales positively with dynamic reasoning questions mixed in.
@@ 5 Discussion
Limitations. We instruction-tune two kinds of MLMs for spatial reasoning, an image-only LLaVA (Liu et al., 2024) and a video LLaVA-Video (Zhang et al., 2024c). While we remember pretraining commonsense as noted in Table 5, we haven’t explored improving other capabilities like math and science reasoning Yue et al. (2024). The scope of the paper, however, is to analyze what kinds of data improve spatial performance, and not a large-scale training of a new MLM. Additionally, further analysis on more MLMs (Dubey et al., 2024; Deitke et al., 2024) might be beneficial.
Future work. Although our study focuses on evaluating the spatial reasoning capabilities of MLMs, it can be extended in various avenues. For instance, to determine the kinds of embodied applications that benefit from improved complex spatial reasoning. We do see gains in accuracy in route planning in VSI-Bench (Yang et al., 2024a), which should help embodied navigation. As a more direct evaluation, we also check the action prediction accuracy (from a choice of going left/right/forward) for a given frame on the SPOC EasyObjectNav benchmark (Ehsani et al., 2024). Our model (SAT Dynamic) scores an accuracy of $51\%$ compared to $40\%$ for training only on static spatial questions. This suggests that embodied navigation might benefit from improved dynamic spatial reasoning. We leave a more thorough evaluation in this direction as future work. Further, leveraging the interactive nature of our scenes with perfect 3D information could facilitate more explorations in dynamic and chain-of-thought causal reasoning.
@@ A.1 Discussion
Limitations. We instruction-tune two kinds of MLMs for spatial reasoning, an image-only LLaVA (Liu et al., 2024) and a video LLaVA-Video (Zhang et al., 2024c). While we remember pretraining commonsense as noted in Table 5, we haven’t explored improving other capabilities like math and science reasoning Yue et al. (2024). The scope of the paper, however, is to analyze what kinds of data improve spatial performance, and not a large-scale training of a new MLM. Additionally, further analysis on more MLMs (Dubey et al., 2024; Deitke et al., 2024) might be beneficial.
Future work. Although our study focuses on evaluating the spatial reasoning capabilities of MLMs, it can be extended in various avenues. For instance, to determine the kinds of embodied applications that benefit from improved complex spatial reasoning. We do see gains in accuracy in route planning in VSI-Bench (Yang et al., 2024a), which should help embodied navigation. As a more direct evaluation, we also checked the action prediction accuracy (from a choice of going left/right/forward) for a given frame on the SPOC EasyObjectNav benchmark (Ehsani et al., 2024). Our model (SAT Dynamic) scores an accuracy of $51\%$ compared to $40\%$ for a model trained only on basic spatial questions. This suggests that embodied navigation might benefit from improved dynamic spatial reasoning. We leave a more thorough evaluation in this direction as future work. Further, leveraging the interactive nature of our scenes with perfect 3D information could facilitate more explorations in dynamic and causal reasoning.
@@ A.2 Tuning details
We tune a widely used open-source MLM, LLaVA-1.5-13B Liu et al. (2024) for fine-tuning experiments. We LoRA-tune (Hu et al., 2021) with rank $256$ and alpha $512$. We full fine-tune the LLaVA-Video-7B model since it is smaller. We see minimal differences in trends by training with more lora parameters or with full fine-tuning as shown in Figure 7. To prevent catastrophic forgetting, we randomly sample examples from the LLaVA Instruct Tuning dataset Liu et al. (2024) with $40\%$ probability while tuning with our synthetic QA pairs. We train all models to around 1 epoch of synthetic data. Due to memory constraints, we use a batch size of $8$ (using gradient accumulation). We set a small learning rate of $5e^{-6}$ (due to our small batch size) with cosine annealing with 1K warm-up steps and a weight decay of 00. Training requires two 48GB NVIDIA GPUs, while inference is possible with one GPU. To demonstrate effect on a video benchmark, we tune another widely used video MLM, LLaVA-Video-7b (Zhang et al., 2024c).

### v3

@@ 3.3 Finetuning MLMs with SAT
We provide the MLM with the required image(s) and the question and ask it to choose between two answers—a correct answer and a distractor answer, following existing benchmark standards (fu2024blink; cai2024temporalbench). The distractor answer is generated by switching the correct answer to the opposite/distractor word (e.g., left $\rightarrow$ right, or “did not move”). To prevent catastrophic forgetting, we mix in some of the original pre-training data during tuning from the LLaVA Instruct-tune dataset (liu2024visual). More tuning and prompt details are in the appendix.
@@ 4 Experiments
Setup. We use two popular open-source MLMs: LLaVA-1.5-13B (liu2024visual) and LLaVA-Video-7B (zhang2024video) for our experiments. We report the standard accuracy metric used for question-answering evaluations by checking if the predicted answer matches the GT answer. We notice better performance for off-the-shelf MLMs when we provide the options in text (e.g., Choose between right or left) as opposed to option numbers (e.g., Choose between option A or B). Hence, we sometimes report higher performances for the baselines than those reported by the original papers. Below, we present the key experimental questions we wish to ask.
@@
CVBench (tong2024cambrian1fullyopenvisioncentric) BLINK (fu2024blink)
@@
Table 4: SAT also improves spatial performance on videos, VSI-Bench (yang2024thinking)
@@
Table 5: SAT also improves out-of domain spatial relation (such as outdoor and video reasoning with humans on MME-RealWorld-Lite) despite having only indoor scenes.
Model Avg Motion Interaction Position Reasoning Perception
LLaVA-Vid-7B 35% 26.8% 24.5% 34% 36.8% 35.0%
+ SAT 39% 33.2% 31.5% 42% 39.7% 39.6%
$\Delta$ Improvement +4% +6.4% +7% +8% +2.9% +4.6%
Table 6: Performance on some traditional spatial benchmarks- VSR (liu2023visualspatialreasoning), GQA-Sp (hudson_gqa_2019)- and other VQA benchmarks- GQA (hudson_gqa_2019), OK-VQA (marino2019okvqavisualquestionanswering), VQAv2 (goyal2017makingvvqamatter)- not focused on spatial relationships. Our spatial tuning remembers pre-training commonsense on other capabilities while improving spatial reasoning.
@@ 4.1 Can SAT improve spatial reasoning in MLMs?
To evaluate the effect of SAT training on spatial performance, we use 4 existing static spatial benchmarks—CVBench (tong2024cambrian1fullyopenvisioncentric), BLINK (spatial subsets) (fu2024blink), Visual Spatial Relations (VSR) (liu2023visualspatialreasoning), and GQA-Spatial (hudson2019gqanewdatasetrealworld). For dynamic reasoning, we use our SAT test sets on both real and synthetic images. Since SAT real is small, we only report the overall average and report individual split performance on our synthetic test set for a more detailed analysis. Finally, we evaluate the effect of SAT tuning on a recent long-video spatial benchmark, VSI-Bench (yang2024thinking), by tuning LLaVA-Video-7B (zhang2024video).
Open MLMs struggle on spatial reasoning. While proprietary MLMs perform better on static reasoning, they still struggle with dynamic reasoning. In Tables 3 and 2, we see that open-source MLMs (rows f and h) struggle on both static relationships and dynamic reasoning. Larger closed-source models like GPT4-o (achiam2023gpt), Gemini-1.5-pro (team2023gemini) and spatially-tuned Robopoint (yuan2024robopointvisionlanguagemodelspatial) are better at static reasoning (in Table 3 row b, d), but still struggle on dynamic QAs (Table 2 rows a-e). On SAT synthetic, aiming at the goal is generally easier for spatially stronger models like RoboPoint and GPT4-o since it also mostly requires judging object position.
Training on SAT significantly improves static spatial reasoning on real images for MLMs. As shown in Table 3, tuning on SAT improves performance on static spatial questions for both LLaVA-1.5 and LLaVA-Video models by significant amounts (rows f vs g and h vs i). Comparing rows a-e, we observe that SAT tuning makes both LLaVA models match/outperform some closed-source models on zero-shot performance on CVBench and BLINK. Compared to an existing spatially-tuned baseline, we outperform RoboPoint-13B (yuan2024robopointvisionlanguagemodelspatial) (row e). This shows promise that SAT instruction-tuning may push performance further for these stronger models.
@@
Camera movement and out-of-domain relations are hard to improve BLINK SpRel has abstract relationships (especially with people) not present in our synthetic data (e.g., “looking away”, “surrounding”). Hence, in Table 3, gains are modest on SpRel (row f vs g), with LLaVA-Video (row h vs i) showing no improvement. Camera movements are also hard (BLINK MV). While more investigation is needed, we believe this may be due to translation invariance in ViT (dosovitskiy2021image) leading to minimal feature changes between images with subtle camera movement.
Training on SAT improves spatial reasoning in videos. Even though SAT only has up to 2 frames in its training data, we evaluate its impact on longer videos on yang2024thinking. The results are shown in Table 4. We see that SAT training helps significantly on route planning capabilities (column Rt. Plan) while also improving other spatial splits. This shows promise that SAT tuning can benefit embodied AI applications.
SAT training also improves on outdoor videos including humans. SAT dataset is rendered using indoor simulators without any humans. However, when trained on SAT, we see improvements on outdoor video reasoning that includes humans on MME-Realworld-Lite (zhang2024mme) as shown in Table 5. While MME-RealWorld has numerous splits related to reasoning and perception, we highlight some of the splits where we see the most improvements when tuned on SAT along with the overall average (Avg column) and the average for all “Reasoning” and “Perception”-based splits.
Our SAT-tuned model maintains its non-spatial reasoning capabilities. We run an evaluation on some standard VQA benchmarks—namely, GQA (hudson2019gqanewdatasetrealworld), VQAv2 (goyal2017makingvvqamatter), and OK-VQA (marino2019okvqavisualquestionanswering) (9K image-QA pairs, random 3K from each). We maintain overall performance on them compared to the off-the-shelf LLaVA (liu2024visual) as shown in Table 6. This suggests that we remember pre-trained vision-language commonsense while adding stronger spatial capabilities.
SAT synthetic is a useful diagnostic test set The lack of improvement on object relations (SpRel and 2DRel) in LLaVA-Video is also reflected by a lesser gain on the GoalAim (row h vs i, Table 2) on SAT synthetic test, where the model needs to judge object position. Similar for SAT EgoM (Table 2) and BLINK MV (Table 3) dealing with camera movements-LLaVA-Video improves more than LlaVA on both. Hence, despite being synthetic, SAT synthetic is also a useful diagnostic test set.
Table 7: Table showing the effectiveness of high-quality but simulated SAT (row d) over real-image but noisier pseudo-annotated spatial data (row b, c). Dynamic spatial reasoning data (row e) further improves static spatial performance.
@@ 4.2 What are effective types of spatial training data?
Sources of instruction tuning Since contemporary work on spatial understanding uses pseudo annotations (chen2024spatialvlm; cheng2024spatialrgptgroundedspatialreasoning), we wish to compare to such baselines as well. Since their datasets are not easily available yet, we reproduce a similar instruction-tuning set using pseudo annotations to create spatial QAs. We keep formats the same as SAT to ensure the performance differences are attributed to the source of spatial information and not the question formats.
  • GQAPseudo: We create SAT-Static like instruction-tuning data by inferring the depth (depthanything) on images from GQA (hudson2019gqanewdatasetrealworld) and VisualGenome (krishna2017visual) and the 2D bounding box annotations. We filter out potential incorrect relationships using simple heuristics (more details in the appendix). We create 225225K static spatial image-QA tuples. Generating pseudo-annotated dynamic QAs is not possible since we cannot easily infer the result of actions on real images, hence, we only generate static QAs.
  • VSR/25VRD: We use spatial relationship datasets like VSR (liu2023visualspatialreasoning) and 2.5VRD (su202125dvisualrelationshipdetection) to produce more real spatial QAs. These datasets contain relationships such as “touching”, or “behind”, allowing us to formulate QAs such as “Is the cat touching the sofa?” We combine VSR and 2.5VRD to create 200200K image-QA tuples.
@@
Figure 4: Ablations with training data size. We see that performance generally scales with data, with dynamic showing better scaling.
SAT simulated data outperforms pseudo-annotations or human annotations. In Table 7, we see that pseudo-annotations can improve performance on certain splits like Count and 2DRel on CVBench (row b). However, noise in annotations (e.g., bounding boxes often not accounting for the entire object) leads to errors in judging depth of the entire object relative to other objects (e.g., a part of the annotated “lake” may be in front of a “tree”, but the majority of it may lie behind it). Following more complex careful curation similar to cheng2024spatialrgptgroundedspatialreasoning may improve this performance, which we leave to future work since their data is not yet easily available. Human-annotated spatial relations (using su202125dvisualrelationshipdetection; liu2023visualspatialreasoning) tend to perform well in-domain on VSR test. While the higher diversity of human-annotated relations also helps prevent forgetting on BLINK SpRel (row c vs d), they are finite and not easily composable to generate varied 3D and dynamic data. Hence, we observe no improvements on SAT dynamic and minimal gains in 3D perception splits.
@@
SAT training scales with more data, especially when dynamic splits added. Since SAT is generated procedurally, we wish to explore whether spatial performance of MLMs scale with more data. Hence, we vary the amounts of training data (train for 1 epoch). As shown in Figure 4, we see that while performance tends to saturate with static-only training (likely due to overfitting to simpler relationships), the performance scales positively with dynamic reasoning questions mixed in.
@@ 5 Discussion
Limitations. We instruction-tune two kinds of MLMs for spatial reasoning, an image-only LLaVA (liu2024visual) and a video LLaVA-Video (zhang2024video). While we remember pretraining commonsense as noted in Table 6, we haven’t explored improving other capabilities like math and science reasoning yue2023mmmu. The scope of the paper, however, is to analyze what kinds of data improve spatial performance, and not a large-scale training of a new MLM. Additionally, further analysis on more MLMs (dubey2024llama; deitke2024molmopixmoopenweights) might be beneficial.
Future work. Although our study focuses on evaluating the spatial reasoning capabilities of MLMs, it can be extended in various avenues. For instance, to determine the kinds of embodied applications that benefit from improved complex spatial reasoning. We do see gains in accuracy in route planning in VSI-Bench (yang2024thinking), which should help embodied navigation. As a more direct evaluation, we also check the action prediction accuracy (from a choice of going left/right/forward) for a given frame on the SPOC EasyObjectNav benchmark (ehsani2024spoc). Our model (SAT Dynamic) scores an accuracy of $51\%$ compared to $40\%$ for training only on static spatial questions. This suggests that embodied navigation might benefit from improved dynamic spatial reasoning. We leave a more thorough evaluation in this direction as future work. Further, leveraging the interactive nature of our scenes with perfect 3D information could facilitate more explorations in dynamic and chain-of-thought causal reasoning. Our causal annotations in language and their corresponding image frames could also advance language-controlled world models genie3, a direction we leave for future work.
@@ A.1 Discussion
Limitations. We instruction-tune two kinds of MLMs for spatial reasoning, an image-only LLaVA (liu2024visual) and a video LLaVA-Video (zhang2024video). While we remember pretraining commonsense as noted in Table 6, we haven’t explored improving other capabilities like math and science reasoning yue2023mmmu. The scope of the paper, however, is to analyze what kinds of data improve spatial performance, and not a large-scale training of a new MLM. Additionally, further analysis on more MLMs (dubey2024llama; deitke2024molmopixmoopenweights) might be beneficial.
Future work. Although our study focuses on evaluating the spatial reasoning capabilities of MLMs, it can be extended in various avenues. For instance, to determine the kinds of embodied applications that benefit from improved complex spatial reasoning. We do see gains in accuracy in route planning in VSI-Bench (yang2024thinking), which should help embodied navigation. As a more direct evaluation, we also checked the action prediction accuracy (from a choice of going left/right/forward) for a given frame on the SPOC EasyObjectNav benchmark (ehsani2024spoc). Our model (SAT Dynamic) scores an accuracy of $51\%$ compared to $40\%$ for a model trained only on basic spatial questions. This suggests that embodied navigation might benefit from improved dynamic spatial reasoning. We leave a more thorough evaluation in this direction as future work. Further, leveraging the interactive nature of our scenes with perfect 3D information could facilitate more explorations in dynamic and causal reasoning.
@@ A.2 Tuning details
We tune a widely used open-source MLM, LLaVA-1.5-13B liu2024visual for fine-tuning experiments. We LoRA-tune (hu2021lora) with rank $256$ and alpha $512$. We full fine-tune the LLaVA-Video-7B model since it is smaller. We see minimal differences in trends by training with more lora parameters or with full fine-tuning as shown in Figure 7. To prevent catastrophic forgetting, we randomly sample examples from the LLaVA Instruct Tuning dataset liu2024visual with $40\%$ probability while tuning with our synthetic QA pairs. We train all models to around 1 epoch of synthetic data. Due to memory constraints, we use a batch size of $8$ (using gradient accumulation). We set a small learning rate of $5e^{-6}$ (due to our small batch size) with cosine annealing with 1K warm-up steps and a weight decay of 0. Training requires two 48GB NVIDIA GPUs, while inference is possible with one GPU. To demonstrate effect on a video benchmark, we tune another widely used video MLM, LLaVA-Video-7b (zhang2024video). FOr LLaVA-Video-7b, the paper results are solely by training on SAT, without mixing any real orginal pretraining data. However, we also do compare the effect of mixing in real pretraining data as show it in Figure 4. We also show the sacling and